OpenStudy (anonymous):

Which expression is equivalent to one over fourn − 16?

OpenStudy (anonymous):

Answer choices?

OpenStudy (anonymous):

1/4 ( n - 4) 1/4 (n- 64) 4 (n-4) 4 (n -64)

OpenStudy (anonymous):

Hmm........ wait a sec plz

OpenStudy (anonymous):

okay. :)

OpenStudy (anonymous):

Well, 1/4 - 16 = negative 15.75

OpenStudy (anonymous):

So, what is an equation that has that same answer? Based on the choices

OpenStudy (misty1212):

lol HI!!

OpenStudy (anonymous):

i think its one of the answers that has a 64

OpenStudy (misty1212):

you really need to use the equation tool

OpenStudy (anonymous):

oh wow lol

OpenStudy (anonymous):

i just need to know the answer.

OpenStudy (anonymous):

Yes, sorry but this is not an answer site. Sorry but my computer malfunctioned

OpenStudy (anonymous):

and i need to learn how to solve the problem

OpenStudy (anonymous):

Ok, first, what is = to n?

OpenStudy (anonymous):

We have to find that out @misty1212 right?

OpenStudy (anonymous):

how do i find out what is equal to n?

OpenStudy (misty1212):

\[4n-16=4(n-4)\]

OpenStudy (misty1212):

that is called factoring

OpenStudy (misty1212):

so \[\frac{1}{4n-16}=\frac{1}{4(n-4)}\]

OpenStudy (anonymous):

okay thankyou very much :)

OpenStudy (anonymous):

Lol you got it?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

:) @misty1212 has the MOST credit :) GOOD LUCK!!!